OpenStudy (anonymous):

Which of the following equations is parallel to the line passing through A(−2, 7), B(−2, −2)? (1 point) y = −2 y = x 3x = 5 −2x = 7y

OpenStudy (anonymous):

y=-2 -2=-2 7=-2 y=x 7=-2 -2=-2 3x=5 3*-2=5 -6=5 3*-2=5 -6=5 -2x=7y -2*-2=7*7 4=49 -2*-2=7*-2 4=-14 none of the choices work

OpenStudy (anonymous):

I didn't think they did either that's why I asked but these are the only choices I have!

OpenStudy (austinl):

You have two points on a line yes? (-2,7) & (-2,-2) To find the slope between the two lines we use, \(\displaystyle slope=\frac{rise}{run}=\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}\) Then we need to look for a line in those given that has the same slope, which means that the two lines are parallel.

OpenStudy (anonymous):

I got -9/4 ..

OpenStudy (austinl):

you should get an x value of 0.

OpenStudy (anonymous):

so y=x is my answer?

OpenStudy (austinl):

I would think so, if you have a slope over zero, that means you have a vertical line.

OpenStudy (anonymous):